OpenStudy (anonymous):

Find f"(x) for each function. Then find f"(0) and f"(2) f(x)=-x/1-x^2

OpenStudy (zzr0ck3r):

\(\frac{-x}{1-x^2}\)?

OpenStudy (anonymous):

Yes

OpenStudy (zzr0ck3r):

do you know how to find the first derivative?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i missed the lesson about this

OpenStudy (zzr0ck3r):

\((\frac{f(x)}{g(x)})'=\frac{f'(x)g(x)-g'(x)f(x)}{(g(x))^2}\)

OpenStudy (zzr0ck3r):

so you have \(f(x) = -x, g(x) = 1-x^2\)

OpenStudy (anonymous):

okay

OpenStudy (zzr0ck3r):

\(f'(x)=-1,g'(x)=-2x \)so \((\frac{-x}{1-x^2})'=\frac{-1(1-x^2)-(-2x)(-x)}{(1-x^2)^2}\)now simplify

OpenStudy (anonymous):

how did 1-x^2 turned to -2x

OpenStudy (anonymous):

i dont know how to simplify this

OpenStudy (zzr0ck3r):

you need to go back and review, it seems you are not at the level you shuold be to do these problems if f(x) = 1-x^2 then f'(x) = -2x

OpenStudy (anonymous):

okay now simplifying this problem will be (-1+1x^2)+(2x+1)/1-x^4

OpenStudy (anonymous):

am i correct

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@@AccessDenied @eyust707

OpenStudy (anonymous):

I need help simplifying this